# Weinstein exactness of nearby Lagrangians and the Lagrangian $C^{0}$ flux conjecture

### v2

@@ 1.3 Applications
For the Hausdorff case, suppose that $r$ is also small enough so that Theorem 3 and Proposition 7 hold in $\mathcal{W}_{r}(L)$. Then, any $L^{\prime}\in\mathcal{L}(L,\tau)$ such that $L^{\prime}\subseteq\mathcal{W}(L)$ must be, up to some global symplectic isotopy, exact in $\mathcal{W}(L)$ — we still denote by $L^{\prime}$ its image under the isotopy. As in the proof of Corollary 17, we note that the path $t\mapsto\Psi(t\Psi^{-1}(L^{\prime}))$, $t\in[0,1]$, is continuous in the Hausdorff metric. Furthermore, it is a Hamiltonian isotopy for all $t>0$. In particular, $L$ must be in the Hausdorff closure of $\mathcal{L}{\mathrm{Ham}}(L^{\prime})\subseteq\mathcal{L}{\mathrm{Symp}_{0}}(L^{\prime})$. But $\mathcal{L}{\mathrm{Symp}_{0}}(L^{\prime})$ is Hausdorff closed by Corollary 8 and the hypotheses on $L$. Therefore, $L^{\prime}\in\mathcal{L}{\mathrm{Symp}_{0}}(L)$.
Putting this result with the Lagrangian flux conjecture, we get the following result.
@@
Let $L$ and $\tau$ be as above. The (path) connected components of $\mathcal{L}(L,\tau)$ in the $C^{1}$ topology are precisely the orbits of $\operatorname{Symp}_{0}(M)$. In particular, the quotient $\mathcal{L}(L,\tau)/\operatorname{Symp}_{0}(M)$ is discrete in the induced topology. If moreover $L$ respects the hypotheses of Theorem 3 or $H_{1}(L;\mathds{R})=0$ and the NLC holds on $T^{*}L$, then the same holds in the Hausdorff metric.
For example, this means that a $\rho$-monotone Clifford torus can never be reached from a Chekanov torus (or any monotone special torus) by a $C^{1}$-continuous path in $\mathcal{L}(\mathds{T}^{2},2\rho)$. Contrast this with the fact that all these tori are Lagrangian isotopic [RGI16].
@@
Combining Corollaries 9 and 18, we get that for all $L\in\mathcal{L}(L,\tau)$, the orbit $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ is both closed and open in $\mathcal{L}(L,\tau)$ with the $C^{1}$ topology. Therefore, $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ must be a union of connected components of $L\in\mathcal{L}(L,\tau)$ by point-set topology. Since $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ is obviously path connected, it must be both a connected component and a path connected component of $L\in\mathcal{L}(L,\tau)$. The proof in the Hausdorff setting is completely analogous.
@@
Going through the proof of Proposition 2 (see Appendix A for details) gives the following quantitative counterpart.
@@
Remarks 4.
There are of course many variations of $c^{e}_{(M,L)}(Q,g)$ that one could take. For example, one could be interested in $c^{A}_{(M,L)}(Q,g)$ or $c^{B}_{(M,L)}(Q,g)$, the largest neighbourhood on which Conjecture A or Conjecture B, respectively, holds. However, if one believes in the NLC, then we should always have $c^{A}=c^{e}$. Moreover, we have not found an example where $c^{B}\neq c^{\mathcal{W}}$. Therefore, $c^{e}$ seems to be the more fruitful version of the relative capacity.
@@ Acknowledgements
We would like to thank Marcelo Atallah for the question that eventually became Conjecture B and explaining, together with Egor Shelukhin, their result for tori [AS24b]. We would also like to extend our gratitude to Frol Zapolsky for indirectly giving the idea for the proof of Proposition 41. Finally, the first and second authors would like to thank Paul Biran and Sobhan Seyfaddini, respectively, for fruitful discussions throughout the writing process.
2 An example to counter them all (but one)
@@
We now explain the construction of the Lagrangian tori from Theorem 1. These tori support the fact that we need in general to require some type of rationality condition on our Lagrangians for Conjectures A, B or C to hold.
@@
We start with the case when $M=\mathds{C}^{3}$. Consider the product torus $L=T(1,2,1+\alpha):=T(1)\times T(2)\times T(1+\alpha)$, where $\alpha>0$ is an irrational number and $T(A)\subseteq\mathds{C}$ denotes the round circle enclosing area $A>0$. By work of Chekanov [Che96], another product torus $T(a,b+a,c+a)$ with $a,b,c>0$ is Hamiltonian isotopic to $L$ in $\mathds{C}^{3}$ if and only if $a=1$ and $\mathrm{span}_{\mathds{Z}}\{b,c\}=\mathrm{span}_{\mathds{Z}}\{1,\alpha\}=:G$.
Since $\alpha$ is irrational, Dirichlet’s approximation theorem ensures that there exist relatively prime integers $p$ and $q$ such that $|p+q\alpha|<\varepsilon$ for any $\varepsilon>0$. This shows that $\mathrm{span}_{\mathds{Z}}\{1,\alpha\}$ is dense in $\mathds{R}$. We now want to show that we may take $b$ and $c$ arbitrarily small and still have them generate $G$. Fix $\varepsilon>0$, and take $b=p+q\alpha$ as above.
Note that $c=r+s\alpha$ will be such that $\mathrm{span}_{\mathds{Z}}\{b,c\}=G$ if and only if the matrix
$\begin{pmatrix}p&r\\ q&s\end{pmatrix}$
has determinant ±$\pm 1$. Pick $r_{0}$ and $s_{0}$ such that the corresponding matrix has determinant ±$\pm 1$ — it exists since $p$ and $q$ are relatively prime. Without loss of generality, we may suppose that $p+q\alpha$ and $r_{0}+s_{0}\alpha$ are positive. Note that if $r$ and $s$ are of the form $r=r_{0}-ip$ and $s=s_{0}-iq$ for some $i\in\mathds{Z}$, then the corresponding matrix also has determinant ±$\pm 1$. But we have that
$|r+s\alpha|=(r_{0}+s_{0}\alpha)-i(p+q\alpha)\qquad\forall i\leq\frac{r_{0}+s_{0}\alpha}{p+q\alpha}.$
Therefore, if we take $i=\lfloor\frac{r_{0}+s_{0}\alpha}{p+q\alpha}\rfloor$, then we have that
$|c|\leq(p+q\alpha)<\varepsilon,$
which proves that $b$ and $c$ may be taken arbitrarily small and still generate $G$.
@@
This means that we can take $T(1,1+b,1+c)$ which are all in the Hamiltonian orbit of $L$ but are arbitrarily $C^{1}$-close to the monotone torus $T(1,1,1)$. Therefore, without the $H$-rational hypothesis on $L$, not even the Lagrangian $C^{1}$ flux conjecture is true in $\mathds{C}^{3}$.
@@
Note that a similar argument as above actually implies that the set of $b,c>0$ such that $T(1,1+b,1+c)$ is Hamiltonian isotopic to $L$ is dense in >$\mathds{R}^{2}_{>0}$. This means that any neighbourhood $U$ of such a torus $T(1,1+b,1+c)$ contains infinitely many $T(1,1+b^{\prime},1+c^{\prime})$ in the same Hamiltonian orbit. But $T(1,1+b,1+c)\cap T(1,1+b^{\prime},1+c^{\prime})=\varnothing$ if $b\neq b^{\prime}$ or $c\neq c^{\prime}$. Therefore, without the $H$-rational hypothesis on $L$, not even Conjecture B is true in $\mathds{C}^{3}$.
@@
We now explain how to generalize the result to any manifold of dimension $2n\geq 6$. First note that by taking a product with $T(1)^{n-3}$, we get a counterexample to our conjectures in $\mathds{C}^{n}$ whenever $n\geq 3$. Furthermore, by Theorem 1.1(ii) of [CS16], the Hamiltonian isotopy from $T(1,\dots,1,1+b,1+c)$ to $T(1,\dots,1,1+b^{\prime},1+c^{\prime})$ can be taken to be fully supported in the ball $B^{2n}(A)$ of capacity $A=n+1+\max\{b+c,b^{\prime}+c^{\prime}\}$, i.e. of radius $\sqrt{\frac{A}{\pi}}$. In particular, for $b$, $b^{\prime}$, $c$, and $c^{\prime}$ small enough, it can be supported in the ball of capacity $n+2$. Therefore, we get a counterexample in $M=B^{2n}(n+2)$ But then, by simply rescaling the ball, we get a counterexample in the ball $B^{2n}(A)$ for any $A>0$. By the Darboux theorem, any symplectic manifold $M^{2n}$ admits a symplectic embedding of the ball $B^{2n}(A)$ for $A$ small enough, which gives the counterexample for every $M$ with $\dim M\geq 6$.
Remarks 5.
@@
Interestingly enough, the above counterexample does not work in dimension 4. Indeed, Chekanov’s classification of product tori implies that every product torus $L$ in $\mathds{C}^{2}$ has a $C^{1}$ neighbourhood $U$ such that $\mathcal{L}{\mathrm{Ham}}(L)\cap U=\{L\}$. In particular, the $C^{1}$ version of Conjecture A holds for $L$, and if its Hamiltonian orbit is not closed, then the limit cannot be a product or a Chekanov torus. By Theorem 1.3 of [CS16], the same holds for product tori in small enough Darboux balls in subtame symplectically aspherical symplectic 4-manifolds.
@@
However, we can use Theorem 1.5 of [CS16] to construct — in a similar fashion as above — a counterexample to Conjecture B in any (spherically) irrational symplectic 4-fold.
@@ 3 Relations between homological rationality and exactness and their standard counterparts
In this section, we discuss relations between standard rationality/exactness and $H$-rationality/$H$-exactness. In Section 3.1, we prove the general following fact: for a closed Lagrangian of the cotangent bundle, being $H$-exact is equivalent to being isotopic to an exact Lagrangian. In Section 3.2, we first explain some specific situations in which $H$-rationality reduces to rationality. We then however give an example which illustrates why we generally work with the $H$-rationality condition rather than the standard rationality one.
@@ 3.1 The central lemma
    $L$ is isotopic to an exact Lagrangian via Lagrangian submanifolds;
@@
  3. (iii)
    $L$ is $H$-exact;
  4. (iv)
    the composition $L\to T^{*}Q\to Q$ is a homotopy equivalence.
@@
Let $i^{*}:H^{1}(T^{*}Q;\mathds{R})\to H^{1}(L;\mathds{R})$ be induced by the inclusion $i:L\to T^{*}Q$. $H$-exactness of $L$ ensures that the form $i^{*}\lambda$ is closed so that it defines a class $[i^{*}\lambda]\in H^{1}(L;\mathds{R})$. Because the canonical projection $\pi:T^{*}Q\to Q$ is a homotopy equivalence, there exists a closed 1-form $\sigma\in\Omega^{1}(Q)$ such that $[i^{*}\lambda]=[i^{*}(\pi^{*}\sigma)]$.
Now, $\sigma$ induces a fibrewise symplectomorphism $\psi_{\sigma}$ of $T^{*}Q$ which satisfies $[\psi_{\sigma}^{*}(i^{*}\lambda)]=0$ so that $\psi_{\sigma}$ maps $L$ to an exact Lagrangian. This shows that (iii) yields (ii), which obviously yields (i).
Note also that (i) implies that the inclusion $L\hookrightarrow T^{*}Q$ is homotopic to the inclusion of an exact Lagrangian. But, when $L$ is exact, the composition $L\to T^{*}Q\to Q$ is a (simple) homotopy equivalence [AK18], i.e. (iv) holds.
Finally, if (iv) holds, then $H_{2}(T^{*}Q,L)=0$, and we have (iii). ∎
3.2 From rationality to $H$-rationality
@@
Here, the rows are the long exact sequences of the pair $(M,L)$ in homotopy and homology with integer coefficients, respectively, and the columns are the various Hurewicz morphisms; it commutes by naturality of the Hurewicz map. We make the abuse of notation of using the same symbols for morphisms in homotopy and homology, since it will be clear from the context which one we are using when.
@@
$h_{1}i(a)=ih^{\prime}_{1}(a)=i\partial(A)=0$
by exactness at $H_{1}(L)$. Therefore, $i(a)\in\operatorname{Ker}h_{1}=[\pi_{1}(M),\pi_{1}(M)]$. By hypothesis, the order $N$ of $[\pi_{1}(M),\pi_{1}(M)]$ is finite, so that $i(Na)=0$. Therefore, there is some $u\in\pi_{2}(M,L)$ such that $\partial(u)=Na$. But note that
@@
By exactness at $H_{2}(M,L)$, there is thus some $B\in H_{2}(M)$ such that $NA=j(B)+h_{2}^{\prime\prime}(u)$. To conclude, we only note that $\omega(j(B))=\omega(B)$ and $\omega(h^{\prime\prime}_{2}(u))=\omega(u)$.
From the above lemma, we directly get the following result.
@@
If $[\pi_{1}(M),\pi_{1}(M)]$ is finite and $\omega(H_{2}(M;\mathds{Z}))\subseteq\omega(\pi_{2}(M))$, then every rational Lagrangian with rationality constant $\tau$ is $H$-rational with rationality constant $\frac{k}{N}\tau$ for some $k\in\mathds{N}$. Moreover, if $\pi_{1}(M)$ is abelian, we have $k=N=1$. In particular, in all those cases, every weakly exact Lagrangian is $H$-exact.
Note that this corollary recovers the statement at the start of the subsection that $H$-rationality and rationality are the same when $\pi_{1}(M)=0$. However, what is perhaps most interesting is the case $M=D^{*}_{r}L$. In that case, the condition on $\omega(H_{2}(M;\mathds{Z}))$ is automatically satisfied, since $\omega$ is exact, and the condition on the commutator subgroup of $\pi_{1}(M)$ becomes that $[\pi_{1}(L),\pi_{1}(L)]$ be finite. We thus get a new version of Theorem 4.
Corollary 27
Let $L$ be a rational Lagrangian submanifold of $(M,\omega)$ such that $[\pi_{1}(L),\pi_{1}(L)]$ is finite, and let $L^{\prime}\in\mathcal{L}{\mathrm{Ham}}(L)$ be a Lagrangian included in a Weinstein neighbourhood $\mathcal{W}_{r}(L)$ of size $r>0$ such that $L^{\prime}$ is weakly exact in $\mathcal{W}_{r}(L)$. Then, for a maybe smaller $r$, $L^{\prime}$ is exact in $\mathcal{W}_{r}(L)$.
Since $H$-exactness implies weak exactness, Theorem 3 also gives a neighbourhood of weak exactness. However, one could also work directly with rationality (see Remark 6 below). Furthermore, there is also a version of Theorem 6 in terms of usual rationality. Therefore, we also have a rational version of the weak Lagrangian $C^{0}$ flux conjecture.
Corollary 28
Let $L$ be a rational Lagrangian in a symplectic manifold $M$ such that $[\pi_{1}(L),\pi_{1}(L)]$ is finite. Suppose that
  1. (i)
    $L$ satisfies the hypotheses of Theorem 3 or $H_{1}(L;\mathds{R})=0$,
  2. (ii)
    and the nearby Lagrangian conjecture holds in $T^{*}L$.
Then, $\mathcal{L}{\mathrm{Ham}}(L)$ and $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ is Hausdorff-closed in $\mathsf{SMan}(L)$.
We end this subsection with an example which showcases the need for $[\pi_{1}(L),\pi_{1}(L)]$ to be finite. This also exemplifies why we are working with $H$-rational Lagrangians and not just rational Lagrangians.
Example.

### v3

@@ 1.3 Applications
For the Hausdorff case, suppose that $r$ is also small enough so that Theorem 4 and Proposition 8 hold in $\mathcal{W}_{r}(L)$. Then, any $L^{\prime}\in\mathcal{L}(L,\tau)$ such that $L^{\prime}\subseteq\mathcal{W}(L)$ must be, up to some global symplectic isotopy, exact in $\mathcal{W}(L)$ — we still denote by $L^{\prime}$ its image under the isotopy. As in the proof of Corollary 17, we note that the path $t\mapsto\Psi(t\Psi^{-1}(L^{\prime}))$, $t\in[0,1]$, is continuous in the Hausdorff topology. Furthermore, it is a Hamiltonian isotopy for all $t>0$. In particular, $L$ must be in the Hausdorff closure of $\mathcal{L}{\mathrm{Ham}}(L^{\prime})\subseteq\mathcal{L}{\mathrm{Symp}_{0}}(L^{\prime})$. But $\mathcal{L}{\mathrm{Symp}_{0}}(L^{\prime})$ is Hausdorff closed by Corollary 9 and the hypotheses on $L$. Therefore, $L^{\prime}\in\mathcal{L}{\mathrm{Symp}_{0}}(L)$.
Together with the Lagrangian flux conjectures, this yields the following result.
@@
Let $L$ and $\tau$ be as above. The (path) connected components of $\mathcal{L}(L,\tau)$ in the $C^{1}$ topology are precisely the orbits of $\operatorname{Symp}_{0}(M)$. In particular, the quotient $\mathcal{L}(L,\tau)/\operatorname{Symp}_{0}(M)$ is discrete in the induced topology. If moreover $L$ respects the hypotheses of Theorem 4 or $H_{1}(L;\mathds{R})=0$ and the NLC holds on $T^{*}L$, then the same holds in the Hausdorff topology.
For example, this means that a $\rho$-monotone Clifford torus can never be reached from a Chekanov torus (or any monotone special torus) by a Hausdorff-continuous path in $\mathcal{L}(\mathds{T}^{2},2\rho)$. Contrast this with the fact that all these tori are Lagrangian isotopic [RGI16].
@@
Combining Corollaries 10 and 18, we get that for all $L\in\mathcal{L}(L,\tau)$, the orbit $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ is both closed and open in $\mathcal{L}(L,\tau)$ with the $C^{1}$ topology. Therefore, $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ must be a union of connected components of $L\in\mathcal{L}(L,\tau)$ by point-set topology. Since $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ is obviously path connected, it must be both a connected component and a path connected component of $L\in\mathcal{L}(L,\tau)$. The proof in the Hausdorff setting is completely analogous.
@@
When $L=S^{1}$, a direct computation gives the following estimate.
@@
Remark 5.
There are of course many variations of $c^{e}_{(M,L)}(Q,g)$ that one could take. For example, one could be interested in $c^{A}_{(M,L)}(Q,g)$ or $c^{B}_{(M,L)}(Q,g)$, the largest neighbourhood on which Speculation A or Speculation B, respectively, holds. However, if one believes in the NLC, then we should always have $c^{A}=c^{e}$. Moreover, we have not found an example where $c^{B}\neq c^{\mathcal{W}}$. Therefore, $c^{e}$ seems to be the more fruitful version of the relative capacity.
@@ Acknowledgements
We thank Georgios Dimitroglou Rizell for asking a question which motivated this project. We would also like to extend our gratitude to Frol Zapolsky for indirectly giving the idea for the proof of Proposition 41. The second author thanks Dylan Cant for pointing out a problem with the original version of Theorem 7. Finally, the second and third authors would like to thank Paul Biran and Sobhan Seyfaddini for fruitful discussions throughout the writing process.
2 Irrational counterexamples
@@
We now explain the construction of the Lagrangian tori from Theorem 1. These tori support the fact that we need in general to require some type of rationality condition on our Lagrangians for Speculations A, B or D to hold.
@@
We start with the case when $M=\mathds{C}^{3}$. Consider the product torus $L=T(1,2,1+\alpha):=T(1)\times T(2)\times T(1+\alpha)$, where $\alpha>0$ is an irrational number and $T(A)\subseteq\mathds{C}$ denotes the round circle enclosing area $A>0$. By work of Chekanov [Che96], another product torus $T(a,a+b,a+c)$ with $a,b,c>0$ is Hamiltonian isotopic to $L$ in $\mathds{C}^{3}$ if and only if $a=1$ and $\mathrm{span}_{\mathds{Z}}\{b,c\}=\mathrm{span}_{\mathds{Z}}\{1,\alpha\}=:G$.
Let $p_{i}/q_{i}$ be the $i$th convergent to $\alpha$ obtained from its infinite continued fraction. In particular $p_{i},q_{i}$ are coprime and $|p_{i}-q_{i}\alpha|<\frac{1}{q_{i+1}}.$ Fix $\varepsilon>0.$ Then for all $i\geq i_{0}$ sufficiently large $|p_{i}-q_{i}\alpha|<\varepsilon.$ Moreover, the matrix
$\begin{pmatrix}p_{i}&q_{i}\\ p_{i+1}&q_{i+1}\end{pmatrix}$
has determinant ±$\pm 1$ and hence for $i\geq i_{0},$ $b=p_{i}-q_{i}\alpha,$ $c=p_{i+1}-q_{i+1}\alpha$ satisfy $\mathrm{span}_{\mathds{Z}}\{b,c\}=G$ while also $|b|<\varepsilon,|c|<\varepsilon.$ By changing the signs of $b,c$ if necessary, which preserves both conditions, we can ensure that $b>0,c>0.$
@@
This means that we can take $T(1,1+b,1+c)$ which are all in the Hamiltonian orbit of $L$ but are arbitrarily $C^{1}$-close to the monotone torus $T(1,1,1)$. This proves (ii) of Theorem 1 in $\mathds{C}^{3}$ and shows that, without the $H$-rational hypothesis on $L$, not even the Lagrangian $C^{1}$ flux conjecture is true in $\mathds{C}^{3}$.
@@
Note that a similar argument as above actually implies that the set of $b,c>0$ such that $T(1,1+b,1+c)$ is Hamiltonian isotopic to $L$ is dense in >$\mathds{R}^{2}_{>0}$. This means that any neighbourhood $U$ of such a torus $T(1,1+b,1+c)$ contains infinitely many $T(1,1+b^{\prime},1+c^{\prime})$ in the same Hamiltonian orbit. But $T(1,1+b,1+c)\cap T(1,1+b^{\prime},1+c^{\prime})=\varnothing$ if $b\neq b^{\prime}$ or $c\neq c^{\prime}$. This proves (i) of Theorem 1 in $\mathds{C}^{3}$ and shows that, without the $H$-rational hypothesis on $L$, not even Speculation B is true in $\mathds{C}^{3}$. Note that this also shows that the Lagrangian flux group of Lagrangian isotopies need not be discrete since that of $T(1,2,1+\alpha)$ is not.
@@
We now explain how to generalize the result to any manifold of dimension $2n\geq 6$. First note that by taking a product with $T(1)^{n-3}$, we get a counterexample to our speculations in $\mathds{C}^{n}$ whenever $n\geq 3$. Furthermore, by Theorem 1.1(ii) of [CS16], the Hamiltonian isotopy from $T(1,\dots,1,1+b,1+c)$ to $T(1,\dots,1,1+b^{\prime},1+c^{\prime})$ can be taken to be fully supported in the ball $B^{2n}(A)$ of capacity $A=n+1+\max\{b+c,b^{\prime}+c^{\prime}\}$, i.e. of radius $\sqrt{\frac{A}{\pi}}$. In particular, for $b$, $b^{\prime}$, $c$, and $c^{\prime}$ small enough, it can be supported in the ball of capacity $n+2$. Therefore, we get a counterexample in $M=B^{2n}(n+2)$ But then, by simply rescaling the ball, we get a counterexample in the ball $B^{2n}(A)$ for any $A>0$. By the Darboux theorem, any symplectic manifold $M^{2n}$ admits a symplectic embedding of the ball $B^{2n}(A)$ for $A$ small enough, which gives the counterexample for every $M$ with $\dim M\geq 6$.
Remark 6.
@@
Interestingly enough, the above counterexample does not work in dimension 4. Indeed, Chekanov’s classification of product tori implies that every product torus $L$ in $\mathds{C}^{2}$ has a $C^{1}$ neighbourhood $U$ such that $\mathcal{L}{\mathrm{Ham}}(L)\cap U=\{L\}$. In particular, the $C^{1}$ version of Speculation A holds for $L$, and if its Hamiltonian orbit is not closed, then the limit cannot be a product or a Chekanov torus. By Theorem 1.3 of [CS16], the same holds for product tori in small enough Darboux balls in subtame symplectically aspherical symplectic 4-manifolds.
@@
However, we expect that one can use Theorem 1.5 of [CS16] to construct — in a similar fashion as above — a counterexample to Speculation B in any (spherically) irrational symplectic 4-fold.
@@ 3 Relations between homological rationality and exactness and their standard counterparts
In this section, we discuss relations between standard rationality/exactness and $H$-rationality/$H$-exactness. In Section 3.1, we prove the general following fact: for a closed Lagrangian of the cotangent bundle, being $H$-exact is equivalent to being isotopic to an exact Lagrangian. In Section 3.2, we explain some specific situations in which $H$-rationality reduces to rationality. Finaly in Section 3.3, we give an example which illustrates why we generally work with the $H$-rationality condition rather than the standard rationality one.
@@ 3.1 The central lemma
    $L$ is $H$-exact;
@@
    the Liouville form $[\lambda_{0}|_{L}]$ is in the image of the map in $H^{1}(\cdot;\mathds{R})$ induced by the composition $L\to T^{*}Q\to Q$;
  3. (iii)
    the composition $L\to T^{*}Q\to Q$ is a homotopy equivalence;
  4. (iv)
    $L$ is isotopic to an exact Lagrangian through Lagrangian submanifolds;
  5. (v)
@@
  6. (vi)
    $L$ is a shift of an exact Lagrangian by a closed one-form on $Q$.
@@
We first show that (i) and (ii) are equivalent. To see this, note that $H$-exactness is equivalent to $\lambda_{0}|_{L}$ vanishes on the image of $H_{2}(T^{*}Q,L;\mathds{R})\to H_{1}(L;\mathds{R})$. But this is equivalent to $\delta^{*}[\lambda_{0}|_{L}]=0$, where $\delta^{*}:H^{1}(L;\mathds{R})\to H^{2}(T^{*}Q,L;\mathds{R})$ is the coboundary operator. By the long exact sequence in cohomology, this is equivalent to $[\lambda_{0}|_{L}]$ being in the image of the map induced by the inclusion $i:L\hookrightarrow T^{*}Q$. We then conclude using the fact that the projection $\pi:T^{*}Q\to Q$ is a homotopy equivalence.
Suppose we have (ii). This means that there exists a closed 1-form $\sigma\in\Omega^{1}(Q)$ such that $[i^{*}\lambda]=[i^{*}(\pi^{*}\sigma)]$. Now, $\sigma$ induces a fibrewise symplectomorphism $\psi_{\sigma}$ of $T^{*}Q$ which satisfies $[\psi_{\sigma}^{*}(i^{*}\lambda)]=0$ so that $\psi_{\sigma}$ maps $L$ to an exact Lagrangian. Furthermore, $\psi_{\sigma}$ generates a shift by $\sigma$. This shows that (ii) yields (vi), which obviously yields (v) and thus (iv).
Note also that (iv) implies by definition that the inclusion $i:L\hookrightarrow T^{*}Q$ is homotopic to the inclusion of an exact Lagrangian. Indeed, $L$ being Lagrangian isotopic to $L^{\prime}$ means precisely that there is a smooth map $F:[0,1]\times L\to T^{*}Q$ such that $F_{0}=i,$ $F_{t}$ for all $t$ is a Lagrangian embedding, and $F_{1}(L)=L^{\prime},$ where $F_{t}:L\to T^{*}Q$ is given by $F_{t}(x)=F(t,x).$ But, when $L$ is exact, the composition $L\to T^{*}Q\to Q$ is a (simple) homotopy equivalence [AK18], i.e. (iii) holds. Finally, note that (iii) directly implies (ii). ∎
Remark 7.
The notions (vi) and (ii) appear in [AS24] under the names of almost exact and $H_{1}$-standard Lagrangians in $T^{*}Q$, respectively.
3.2 When does $H$-rationality reduce to rationality?
@@
Here, the rows are the long exact sequences of the pair $(M,L)$ in homotopy and homology with integer coefficients, respectively, and the columns are the various Hurewicz morphisms; it commutes by naturality of the Hurewicz map.
@@
$h_{1}i(a)=ih^{\prime}_{1}(a)=i\partial(A)=0.$
Therefore, $i(a)\in\operatorname{Ker}h_{1}=[\pi_{1}(M),\pi_{1}(M)]$. By hypothesis, $i(Na)=0$. Therefore, there is some $u\in\pi_{2}(M,L)$ such that $\partial(u)=Na$. But note that
@@
There is thus some $B\in H_{2}(M)$ such that $NA=j(B)+h_{2}^{\prime\prime}(u)$. To conclude, we only note that $\omega(j(B))=\omega(B)$ and $\omega(h^{\prime\prime}_{2}(u))=\omega(u)$.
Note that this corollary recovers the statement at the start of the subsection that $H$-rationality and rationality are the same when $\pi_{1}(M)=0$. In the case $M=D^{*}_{r}L$, the condition that $\omega(H_{2}(M;\mathds{Z}))=0$ is automatically satisfied since $\omega$ is exact, and the condition on the commutator subgroup of $\pi_{1}(M)$ becomes that $[\pi_{1}(L),\pi_{1}(L)]$ be finite. Therefore, we get directly the following result.
@@
Suppose that $[\pi_{1}(L),\pi_{1}(L)]$ is finite. Then, Conjecture C is equivalent to its homotopical version, i.e. the one where $K$ is only rational, and we get a bound on its rationality constant.
Remark 8.
The homotopical version of Theorem 2, that is, Theorem 39 below, proves the homotopical version of Conjecture C when $L=L_{0}\times\mathds{T}^{m}$ with $H_{1}(L_{0};\mathds{R})=0$. This intersects with the above corollary precisely in the case when $\pi_{1}(L_{0})$ is finite.
Using Corollary 26, we can obtain a homotopical version of Theorem 4. However, the proof of Theorem 5 truly requires $H$-exactness. One exception to this is when $\pi_{2}(M,L)\to\pi_{1}(L)$ is surjective, but an argument similar to the above shows that rationality is then equivalent to $H$-rationality as long as $q\omega(H_{2}(M;\mathds{Z}))=\omega(\pi_{2}(M,L))$ for some $q\in\mathds{Q}$, which is the case generically (and is always the case if $[\pi_{1}(L),\pi_{1}(L)]$ is finite).
3.3 Why $H$-rationality in general?
We end this section with an example which showcases the need to work with $H$-rational Lagrangians and not just rational Lagrangians for many applications. Note that this example is such that $[\pi_{1}(L),\pi_{1}(L)]$ is not finite.